# Bangkok 10330, Thailand

### v1

@@ 1 Introduction
$V(R)={\frac{M_{p}^{2}}{M_{*}^{3}}}{\frac{\Lambda_{5}\ \langle R\rangle}{R}}\,,$ (4)
@@
$V(R)=\frac{2\pi\ \Lambda_{5}\ \langle R\rangle^{2}}{R}+\left(\frac{\langle R\rangle}{R}\right)^{2}\ T_{4}+V_{C}(R)\,,$ (6)
@@
stands for the quantum corrections to the vacuum energy due to Casimir forces, with $m_{i}=R_{i}^{-1}$ the masses of the 5-dim fields, $\Theta$ a step function, and $N_{F}-N_{B}$ the difference between the number of light fermionic and bosonic degrees of freedom. The value and in particular the sign of the the uplifted vacuum energy $V(R)$ now depends on the difference $N_{F}-N_{B}$ Anchordoqui:2023woo .
@@
A minimal set up that accommodates the AdS $\to$ dS transition requires a 5-dim mass spectrum containing: the graviton, three generations of light right-handed neutrinos, and a light real scalar field $\phi$. In Anchordoqui:2023woo it was hypothesised that the real scalar has a potential $V(\phi)$ holding two local minima with very small difference in vacuum energy and bigger curvature (mass) of the lower one, such that at $z_{c}\sim 2$ the false vacuum could “tunnel” to its true vacuum state. After the quantum tunneling $\phi$ becomes more massive and its contribution to the Casimir energy becomes exponentially suppressed. The tunneling process then changes the difference between the total number of fermionic $N_{F}$ and bosonic $N_{B}$ degrees of freedom contributing to the quantum corrections of the vacuum energy. As shown in Anchordoqui:2023woo the change of $N_{F}-N_{B}$ could drive the required AdS $\to$ dS transition at $z_{c}\sim 2$ if $\Lambda_{5}^{1/5}\sim 22.6~{}{\rm meV}$, $|T_{4}|^{1/4}\sim 24.2~{}{\rm meV}$, and $\langle R\rangle\sim 10~{}\mu{\rm m}$. In this paper we particularise the study carried out in Antoniadis:2024ent to determine the shape of the potential $V(\phi)$ that would allow $\phi$ to undergo the hypothesised Coleman-de Luccia (CdL) transition Coleman:1980aw and at the same time would exponentially suppress the contribution of the scalar field to $V_{C}(R)$ for $z\lesssim 2$.
The layout of the paper is as follows. In Sec. 2, we briefly review the basics of the bounce configuration and their relation to the second functional variation of the Euclidean action, which has a negative mode Callan:1977pt . The existence and uniqueness of such a negative mode is a criterion for the bounce to mediate vacuum transitions. We thus investigate the interplay of negative and Kaluza-Klein (KK) modes. We demand that the system has a single negative mode and therefore a KK tower with positive modes. This condition leads to a stability constraint on the radius of the compact dimension. In Secs. 3 and 4, we set up conditions on the shape of the potential $V(\phi)$ under which a late time AdS $\to$ dS transition could be driven by Casimir forces. In Sec. 5 we equate the required lifetime of the false vacuum to the CdL (semi-classical) tunneling rate per unit volume to obtain an estimate of the on-shell Euclidean action of the bounce. The normalization factor, which is a one-loop determinant, is approximated using dimensional analysis. Because of the exponential dependence of the vacuum decay rate on the bounce factor, this approximation does not lead to large errors in our calculations. Adjusting the free parameters of the model to fiducial values we study two possible regimes of the bounce factor. We show that to accommodate the required transition at $z_{c}$ the size of the CdL instanton must be larger than the compactification radius. The paper wraps up in Sec. 6 with some conclusions.
@@ 2.1 Non-compact $d$-dimensions
This is not in the form of the eigenmode equation (14) because its right hand side depends on $\xi$. However, in the thin-wall approximation Coleman:1977py ; Coleman:1980aw , the bounce solution exhibits non-trivial dependence on $\xi$ only on the wall, which is located at $\xi=\bar{\xi}$, having a constant profile in- and outide the wall. This implies that $\dot{\phi}_{\mathrm{B}}$ has a sharp peak around the wall $\xi=\bar{\xi}$ and is zero in- and outside the wall. Therefore, Eq. (21) can be rewritten as
@@ 2.2 $d$-dimensions with a compact direction
where we used the decomposition of the Laplacian (15) with the replacement $d\to d-1$. Since $\phi_{\mathrm{B}}$ is independent of $y$, this has a separable form. To find mode functions for fluctuations, we need to specify boundary conditions on $\eta$ at the two endpoints $y=0,\pi R_{0}$ of the interval $S^{1}/\mathbf{Z}_{2}$. In general, $\eta$ can be expanded in $\sin$ and $\cos$ wave functions. Since fluctuations without extra-dimensional dependence are allowed to contribute the one-loop determinant, we need a zero (constant) mode. We therefore adopt $\cos$ wavefunctions satisfying the Neumann boundary conditions. This choice is consistent with the classical bounce configuration since it is independent of $y$ and hence satisfies the Neumann boundary condition trivially. Taking them into account, we can express each eigenfunction in the form:
@@
In summary, the set of all eigenvalues of the eigenmode equation (29) is obtained as follows. We expand the set of all eigenvalues of $\tilde{O}_{{\mathrm{B}},\ell}$ to the set ${\sigma}_{\mathrm{KK}}(\tilde{O}_{{\mathrm{B}},\ell})$ by assigning the KK tower of the form $\{h_{\ell}+s^{2}/R_{0}^{2}\}_{s=0,1,2,\cdots}$ to each eigenvalue $h_{\ell}$ of $\tilde{O}_{{\mathrm{B}},\ell}$, and then collect the sets ${\sigma}_{\mathrm{KK}}(\tilde{O}_{{\mathrm{B}},\ell})$ for all spins $\ell$; namely $\cup_{\ell\geq 0}\,{\sigma}_{\mathrm{KK}}(\tilde{O}_{{\mathrm{B}},\ell})$. Each eigenvalue in ${\sigma}_{\mathrm{KK}}(\tilde{O}_{{\mathrm{B}},\ell})$ corresponds to the eigenfunctions of the form in (30) and has multiplicity $\mathfrak{d}_{\ell,d-1}$.
@@ 3 Incompatibility of quartic potentials with hypotheses of the transition
We start from the requirement that the quartic scalar potential $V(\phi)$ should have two minima at $v_{+}$ and $v_{-}$. Using the degrees of freedom of shifting $\phi$ and adding a constant to the potential, we are left with three real parameters out of five in a quartic potential. The requirement for the two minima is then sufficient to fix $V^{\prime}(\phi)$ to
@@
As we already mentioned in Section 2, we adopt the thin-wall approximation Coleman:1977py ; Coleman:1980aw , which is justified when the energy difference ${\varepsilon}$ is sufficiently small, making the thickness of the bounce much smaller than the size of the wall. This condition is translated into our parameters as
@@ 4 Sixth order potential
so that the difference of the vacuum energies is still ${\varepsilon}$ [^3] Exactly speaking, the vacuum expectation values of $\phi$ at the two vacua $\hat{v}_{\pm}$ of the tunneling potential $V$ are different from $v_{\pm}$: $\hat{v}_{+}=v_{+}+\mathcal{O}({\varepsilon}^{1/3})$ and $\hat{v}_{-}=v_{-}+\mathcal{O}({\varepsilon})$. The vacuum energies then change into $V(\hat{v}_{+})={\varepsilon}+\mathcal{O}({\varepsilon}^{4/3})$ and $V(\hat{v}_{-})=\mathcal{O}({\varepsilon}^{2})$. Therefore, the difference of the vacuum energies is still ${\varepsilon}+\mathcal{O}({\varepsilon}^{4/3})$. A schematic representation of the potential is shown in Fig. 2.
@@ 5 Bounce factor from dimensional analysis and lifetime prerequisite
where $\hat{\kappa}_{5}:=\kappa_{5}S_{1}^{2}/{\varepsilon}$ is a dimensionless version of the five-dimensional reduced Planck length to the third power $\kappa_{5}$ introduced in Antoniadis:2024ent , controlling the gravitational correction to the bounce in the presence of gravity. Next, we consider the gravity-decoupling limit, in which only the overall coefficients matter. In the thin-wall approximation, the scalar field $\phi$ varies on the wall, while the dimensionless scale factor $\rho$ of the spherical bubble (with radius $H^{-1}$) is considered nearly constant $\bar{\rho}$. Then, the result for the compact dimension should be valid in the regime where the radius $R_{0}$ of the compact dimension is smaller than the radius $\bar{\rho}H^{-1}\simeq S_{1}/\varepsilon$ of the 3-sphere. When $R_{0}$ is larger than the latter, one should expect to reproduce the 5-dim non-compact result for the bounce, which is achieved by letting $R_{0}\simeq S_{1}/\varepsilon$. This is quantified in the negative mode analysis of the CdL instanton discussed in Sec. 2. Substituting
@@
Now, (76) yields ${\delta}\lesssim 0.01$ (see Eq. (63)). The lifetime of the false vacuum is related to the decay rate (8) by
@@
The primed determinant $\det^{\prime}O_{\text{B}}$ is the determinant of $O_{\text{B}}$ with four zero modes with spin 1 removed, and $\det O_{\text{FV}}$ is the full determinant of $O_{\text{FV}}$. Each zero mode of $O_{\mathrm{B}}$ corresponds to each modulus of the bounce and hence the number of the zero modes is $D$. Since $O_{\mathrm{B}}$ and $O_{\mathrm{FV}}$ have mass dimension 2, the ratio ${\det^{\prime}O_{\text{B}}}/{\det O_{\text{FV}}}$ has dimension $-2D$ and hence $A$ has dimension $D$. See Appendix A.1 for more details.
@@
For concreteness, we assume $R_{0}\sim 10\,\mu\text{m}\sim 50\,\text{eV}^{-1}$ and $\varepsilon\lesssim(2\,\mbox{meV})^{5}$, yielding
@@
    $\frac{S_{1}}{\varepsilon}<\frac{R_{0}}{\sqrt{3}}\sim 29~{}\text{eV}^{-1}.$ (91)
@@
    $\frac{S_{1}}{\varepsilon}>\frac{R_{0}}{\sqrt{3}}\sim 29~{}\text{eV}^{-1}.$ (93)
@@
where we used $R_{0}\sim 50\,\text{eV}$ and $M_{*}\sim 10^{9}\,\text{GeV}$. However, this big lower bound means the breakdown of the effective theoretical treatment of our 5-dim action of $\phi$, and it is also impossible in the context of our analysis due to the upper bound on ${\varepsilon}$ given by (77).

### v2

@@ 1 Introduction
$V_{\Lambda_{5}}(R)={\frac{M_{p}^{2}}{M_{*}^{3}}}{\frac{\Lambda_{5}\ \langle R\rangle}{R}}\,,$ (4)
@@
$V_{\mathrm{eff}}(R)=\frac{2\pi\ \Lambda_{5}\ \langle R\rangle^{2}}{R}+\left(\frac{\langle R\rangle}{R}\right)^{2}\ T_{4}+V_{C}(R)\,,$ (6)
@@
stands for the quantum corrections to the vacuum energy due to Casimir forces, with $m_{i}=R_{i}^{-1}$ the masses of the 5-dim fields, $\Theta$ a step function, and $N_{F}-N_{B}$ the difference between the number of light fermionic and bosonic degrees of freedom. The value and in particular the sign of the the uplifted vacuum energy $V_{\mathrm{eff}}(R)$ now depends on the difference $N_{F}-N_{B}$ Anchordoqui:2023woo .
@@
A minimal set up that accommodates the AdS $\to$ dS transition requires a 5-dim mass spectrum containing: the graviton, three generations of light right-handed neutrinos, and a light real scalar field $\phi$. In Anchordoqui:2023woo it was hypothesised that the real scalar has a potential $V(\phi)$ holding two local minima with very small difference in vacuum energy and bigger curvature (mass) of the lower one, such that at $z_{c}\sim 2$ the false vacuum could “tunnel” to its true vacuum state. After the quantum tunneling $\phi$ becomes more massive and its contribution to the Casimir energy becomes exponentially suppressed. The tunneling process then changes the difference between the total number of fermionic $N_{F}$ and bosonic $N_{B}$ degrees of freedom contributing to the quantum corrections of the vacuum energy. As shown in Fig. 1, the change of $N_{F}-N_{B}$ could drive the required AdS $\to$ dS transition at $z_{c}\sim 2$ (keeping the size of the compact space approximately fixed), if $\Lambda_{5}^{1/5}\sim 22.6~{}{\rm meV}$, $|T_{4}|^{1/4}\sim 24.2~{}{\rm meV}$, and $\langle R\rangle\sim 10~{}\mu{\rm m}$ [^1] We take $h=c=1$ and so $1~{}\mu{\rm m}\simeq 0.806~{}{\rm eV}^{-1}$.
Figure 1: The potential $V_{\mathrm{eff}}(R)$ for $(\Lambda_{5})^{1/5}=22.6~{}{\rm meV}$ and $|T_{4}|^{1/4}=24.2~{}{\rm meV}$, considering $N_{F}-N_{B}=6$ (AdS) and $N_{F}-N_{B}=7$ (dS). From Ref. Anchordoqui:2023woo .
More concretely, the compactification radius changes from $R_{0}\simeq 9.0~{}\mu{\rm m}$ (in the AdS phase) to $R_{0}\simeq 9.4~{}\mu{\rm m}$ (in the dS phase). This corresponds to a variation in Newton’s gravitational constant of
$\frac{\Delta G_{N}}{G_{N}}=\frac{\Delta R_{0}}{R_{0}}\simeq 0.04\,.$ (8)
A plethora of investigations have set constraints on the variation of $\Delta G_{N}/G_{N}$ over cosmic timescales. An early study, based on the abundances of light elements (i.e. D, 3He, 4He, and 7Li) yielded $-0.3<\Delta G_{N}/G_{N}<0.4$ Accetta:1990au . This study was carried out on the assumption that other physical constants remained fixed. Subsequently, a study which incorporated constraints from extra relativistic degrees of freedom translated the limits on the speed-up factor of the early universe into bounds on $\Delta G_{N}/G_{N}$, leading to $-0.10<\Delta G/G_{N}<0.13$ Cyburt:2004yc . More recent work refined this range to $-0.040<\Delta G_{N}/G_{N}<0.006$, using improved observational data and updated cosmological assumptions Yeh:2022heq . It is important to stress that the bound reported in Yeh:2022heq follows from $\Delta G_{N}/G_{N}=(7/43)(N_{\rm eff}-3)$, using an effective number of neutrino-like species $N_{\rm eff}=2.898\pm 0.141$. Should we use instead the 95%CL bound reported by the Planck Collaboration $N_{\rm eff}<3.34$ Planck:2018vyg , the upper limit on $\Delta G_{N}/G_{N}<0.05$ is consistent with (8). Using the PRIMAT code and assuming variations only in $G$ while keeping all other cosmological parameters fixed leads to: $\Delta G_{N}/G_{N}=0.01^{+0.06}_{-0.05}$ Alvey:2019ctk . In addition, WMAP and SDSS data constrain the total variation of $\Delta G_{N}/G_{N}$ since the epoch of recombination as $-0.083<\Delta G_{N}/G_{N}<0.095$ at $2\sigma$CL Wu:2009zb . Last but not least, combining information from Planck 2018 CMB temperature, polarization, and lensing, together with a compilation of BAO measurements from the Baryon Oscillation Spectroscopic Survey (BOSS), leads to $-0.13<\Delta G_{N}/G_{N}<0.049$ at $2\sigma$CL Ballardini:2021evv . Thus, we conclude that the change in radius is small enough so that there is no conflict with observations. Actually, the variation of $\Delta G_{N}/G_{N}$ is within the observationally motivated range (estimated in Efstratiou:2025xou to be between 1% and 7%).
[…]
@@ 2.1 Non-compact $d$-dimensions
This is not in the form of the eigenmode equation (15) because its right hand side depends on $\xi$. However, in the thin-wall approximation Coleman:1977py ; Coleman:1980aw , the bounce solution exhibits non-trivial dependence on $\xi$ only on the wall, which is located at $\xi=\bar{\xi}$, having a constant profile in- and outside the wall. [^4] See also Appendix B for the thin-wall approximation. This implies that $\dot{\phi}_{\mathrm{B}}$ has a sharp peak around the wall $\xi=\bar{\xi}$ and is zero in- and outside the wall. Therefore, Eq. (22) can be rewritten as
@@ 2.2 $d$-dimensions with a compact direction
where we used the decomposition of the Laplacian (16) with the replacement $d\to d-1$. Since $\phi_{\mathrm{B}}$ is independent of $y$, this has a separable form. To find mode functions for fluctuations, we need to specify boundary conditions on $\eta$ at the two endpoints $y=0,\pi R_{0}$ of the interval $S^{1}/\mathbf{Z}_{2}$. In general, $\eta$ can be expanded in $\sin$ and $\cos$ wave functions. Since fluctuations without extra-dimensional dependence are allowed to contribute to the one-loop determinant, we need a KK zero (constant) mode. We therefore adopt $\cos$ wavefunctions satisfying the Neumann boundary conditions. This choice is consistent with the classical bounce configuration since it is independent of $y$ and hence satisfies the Neumann boundary condition trivially. Taking them into account, we can express each eigenfunction in the form:
@@
In summary, the set of all eigenvalues of the eigenmode equation (30) is obtained as follows. We extend the set of all eigenvalues of $\tilde{O}_{{\mathrm{B}},\ell}$ to the set ${\sigma}_{\mathrm{KK}}(\tilde{O}_{{\mathrm{B}},\ell})$ by assigning the KK tower of the form $\{h_{\ell}+s^{2}/R_{0}^{2}\}_{s=0,1,2,\cdots}$ to each eigenvalue $h_{\ell}$ of $\tilde{O}_{{\mathrm{B}},\ell}$, and then collect the sets ${\sigma}_{\mathrm{KK}}(\tilde{O}_{{\mathrm{B}},\ell})$ for all spins $\ell$; namely $\cup_{\ell\geq 0}\,{\sigma}_{\mathrm{KK}}(\tilde{O}_{{\mathrm{B}},\ell})$. Each eigenvalue in ${\sigma}_{\mathrm{KK}}(\tilde{O}_{{\mathrm{B}},\ell})$ corresponds to the eigenfunctions of the form in (31) and has multiplicity $\mathfrak{d}_{\ell,d-1}$.
@@ 3 Incompatibility of quartic potentials with hypotheses of the transition
We start from the requirement that the quartic scalar potential $V(\phi)$ in 5-dim should have two minima at $v_{+}$ and $v_{-}$. Using the degrees of freedom of shifting $\phi$ and adding a constant to the potential, we are left with three real parameters out of five in a quartic potential. The requirement for the two minima is then sufficient to fix $V^{\prime}(\phi)$ to
@@
As we already mentioned in Section 2, we adopt the thin-wall approximation Coleman:1977py ; Coleman:1980aw , which is justified when the energy difference ${\varepsilon}$ is sufficiently small, making the thickness of the bounce much smaller than the size of the wall (see Appendix B). This condition is translated into our parameters as
@@ 4 Sixth order potential
so that the difference of the vacuum energies is still ${\varepsilon}$ [^6] Exactly speaking, the vacuum expectation values of $\phi$ at the two vacua $\hat{v}_{\pm}$ of the tunneling potential $V$ are different from $v_{\pm}$: $\hat{v}_{+}=v_{+}+\mathcal{O}({\varepsilon}^{1/3})$ and $\hat{v}_{-}=v_{-}+\mathcal{O}({\varepsilon})$. The vacuum energies then change into $V(\hat{v}_{+})={\varepsilon}+\mathcal{O}({\varepsilon}^{4/3})$ and $V(\hat{v}_{-})=\mathcal{O}({\varepsilon}^{2})$. Therefore, the difference of the vacuum energies is ${\varepsilon}+\mathcal{O}({\varepsilon}^{4/3})$. A schematic representation of the potential is shown in Fig. 3.777In this figure, the mass ratio $m_{-}/m_{+}$ was set to a number much larger than 1 for visualisation. However, as we will see in the next section, it is sufficient to assume $m_{-}/m_{+}\gtrsim 5$ for the AdS $\to$ dS transition.
@@ 5 Bounce factor from dimensional analysis and lifetime prerequisite
where $\hat{\kappa}_{5}:=\kappa_{5}S_{1}^{2}/{\varepsilon}$ is a dimensionless version of the five-dimensional reduced Planck length to the third power $\kappa_{5}$ introduced in Antoniadis:2024ent , controlling the gravitational correction to the bounce in the presence of gravity. Next, we consider the gravity-decoupling limit, in which only the overall coefficients matter. In the thin-wall approximation, [^8] We verify in Appendix B that the choice of the parameters discussed later in this section is consistent with the thin-wall approximation. the scalar field $\phi$ varies on the wall, while the dimensionless scale factor $\rho$ of the spherical bubble (with radius $H^{-1}$) is considered nearly constant $\bar{\rho}$. Then, the result for the compact dimension should be valid in the regime where the radius $R_{0}$ of the compact dimension is smaller than the radius $\bar{\rho}H^{-1}\simeq S_{1}/\varepsilon$ of the 3-sphere. When $R_{0}$ is larger than the latter, one should expect to reproduce the 5-dim non-compact result for the bounce, which is achieved by letting $R_{0}\simeq S_{1}/\varepsilon$. This is quantified in the negative mode analysis of the CdL instanton discussed in Sec. 2. Substituting
@@
Before proceeding, we pause to clarify that even though for $R_{0}<\sqrt{3}\ S_{1}/\varepsilon$, the decay rate is described by an effective 4D instanton, this decay rate describes transitions between 5D vacua. Of critical interest for our investigation is that in the intial vacuum the mass of the scalar field is smaller than the compactification radius, but after the transition the mass is bigger than the compactification radius. Therefore, the transition should not be analyzed purely from a 4D perspective; namely, the field should be treated as 5D or else we must keep its KK excitations.
@@
Moreover, $m_{+}$ should be much smaller than the compactification scale $R_{0}^{-1}$ and bigger than the 5D cosmological constant $\Lambda_{5}^{1/5}$, which fixes its value around $m_{+}\sim 100$ meV. Now, (77) yields ${\delta}\lesssim 0.01$ (see Eq. (64)). The lifetime of the false vacuum is related to the decay rate (9) by
@@
The primed determinant $\det^{\prime}O_{\text{B}}$ is the determinant of $O_{\text{B}}$ with $D$ zero modes of spin 1 removed, and $\det O_{\text{FV}}$ is the full determinant of $O_{\text{FV}}$. Recall that each zero mode of $O_{\mathrm{B}}$ corresponds to each modulus of the bounce and hence the number of the zero modes is $D$. Since $O_{\mathrm{B}}$ and $O_{\mathrm{FV}}$ have mass dimension 2, the ratio ${\det^{\prime}O_{\text{B}}}/{\det O_{\text{FV}}}$ has dimension $-2D$ and hence $A$ has dimension $D$. See Appendix A.1 for more details.
@@
For concreteness, we assume $R_{0}\sim 10\,\mu\text{m}\sim 8\,\text{eV}^{-1}$ and $\varepsilon\lesssim(2\,\mbox{meV})^{5}$, yielding
@@
    $\frac{S_{1}}{\varepsilon}<\frac{R_{0}}{\sqrt{3}}\sim 6~{}\text{eV}^{-1}.$ (92)
@@
    $\frac{S_{1}}{\varepsilon}>\frac{R_{0}}{\sqrt{3}}\sim 6~{}\text{eV}^{-1}.$ (94)
@@
Armed with our finding we now provide an assessment on the size of the bubble and the possiblity of having bubble collisions. The size of the bubble is estimated in Eq. (74) and is given by $\bar{\xi}=3S_{1}/\varepsilon$. We demand that $\bar{\xi}$ is bigger than a minimum length which is proportional to the compactification scale to avoid the negative mode instability. However, we stress that this is a lower bound for the size of the bubble. To find an upper bound, we first note that once the lifetime of the 5D false vacuum $\tau$ and the true-vacuum mass $m_{-}$ are fixed (see Eqs. (76) and (85)), the value of $B$ is essentially determined (see Eq. (86)). Now, we can conclude that $S_{1}/\varepsilon$ as given by Eq. (88) corresponds to the probability of a decay at $z=2$. As previously noted, Eq. (93) provides a lower bound for $S_{1}/\varepsilon$, which is $720~{}{\rm eV}^{-1}$. But of course we can increase the size of the bubbles by decreasing $\varepsilon$ while keeping fixed the probability determined by $B$. From Eq. (88) this means that $S_{1}/\varepsilon\propto\varepsilon^{-1/4}$. This in turn means that parametrically we can make the size of the bubble as large as the size of the universe. Obviously, the limit must be taken with $S_{1}$ and $\varepsilon$ going to zero with different rates and this is dictated by the condition that $B$ remains constant.
In summary, the size of the bubble is a parameter that can be assessed in our model. The size is not fixed, but for fixed probability it depends parametrically on $\varepsilon$. The size then can be arbitrarily large parametrically, and as a matter of fact it can be as large as the size of the universe at the time of the transition.
In practice, if the bubble is very large we expect no bubble collisions. As we start decreasing the size, the dynamics becomes more and more important because collisions will start playing some role. Actually, there may be a critical size for which the process destabilizes. Indeed, if the bubble size $\bar{\xi}$ is much smaller than the horizon scale $\sim 1/H_{0}$, one may need a dedicated analysis of dynamics of the bubbles, which is beyond the scope of this work.
@@
where we used $R_{0}\sim 8\,\text{eV}$ and $M_{*}\sim 10^{9}\,\text{GeV}$. However, this big lower bound means the breakdown of the effective theoretical treatment of our 5-dim action of $\phi$, and it is also impossible in the context of our analysis due to the upper bound on ${\varepsilon}$ given by (78).